# Automatic Input Feature Relevance via Spectral Neural Networks

### v2

@@ 3.2 Correlated Gaussian distributions dataset
$\begin{split}&x_{2n}\sim\mathcal{N}(0,1)\\ &x_{2n+1}=x_{2n}\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\text{if}\;\;c(\vec{x})=0\\ &x_{2n+1}=\begin{cases}x_{2n}\;\text{with prob.}\;1-p_{n}\\ -x_{2n}\;\text{with prob.}\;p_{n}\end{cases}\;\;\text{if}\;\;c(\vec{x})=1\end{split}$ (2)
where $c(\vec{x})$ is the target class of the input and the index $n\in\{0,1,2,3\}$ denotes the four pairs. The last two components are independently drawn from two identical normal distributions $\mathcal{N}(1,0)$ regardless of the class of belonging:
@@
To better understand the relationship between the different components, let us focus on the first two entries of the input vectors, $x_{0}$ and $x_{1}$. For data belonging to class 00 (as we label the first of the two) $x_{0}=x_{1}$, that is, the first two components are perfectly correlated. Conversely, the two components of data belonging to the other class (denoted by $1$) are equal with probability $1-p_{0}$ and have opposite sign with probability $p_{0}$. If $p_{0}=0$, the first two components are equal for both classes. No information about the class to which the supplied data belong can be thus extracted by comparing the two above components. The second component is indeed a mere copy of the first and therefore irrelevant for classification purposes. On the other hand, if $p_{0}=1$, the first and second components from class $1$ data are in perfect anti-correlation, while data from class 00 are still matched and thus fully correlated. Hence, in the latter case, it is possible to completely separate the two classes by solely comparing $x_{0}$ and $x_{1}$. The Parameter $p_{n}$ quantifies therefore the degree of sensible information which can be pulled out from a direct comparison of the two components in the $n$-th pair. We want to stress that, regardless of the value of $p_{n}$, neither component individually contains information on the specific class to which the vector belongs. Therefore, useful information can be obtained to solve the classification task only by analyzing the two components simultaneously. Following this observations, we generated the dataset by using four different values of $p_{n}\in[0,1]$ for the four pairs. Finally, the last two components are not correlated with the others and are identically distributed for both classes, hence void of information. We therefore assign a parameter $p_{4}=0$ to these two components to explicitly state their manifest irrelevance.
@@
The architecture of the employed neural network, parameterized via its spectral modality, is identical to that put forward for the analysis of the uncorrelated dataset (except for the first layer that in this case receives just 10 entries). In upper right panel of Fig. 4, the post-training eigenvalues referred to the input layer are shown. Firstly, we notice that pairs are correctly spotted out, as they are assigned the very same eigenvalue entry. Moreover, eigenvalues associated with irrelevant components are tuned zero for this dataset as well, an observation that testifies again on the ability of the spectral network to discriminate between irrelevant and relevant features. Lastly, the magnitude of the recorded eigenvalues scales proportionally to the control parameter $p_{n}$, which quantifies the relevance of the selected pair for classification purposes (bottom right panel, Fig. 4). In conclusion, and also for this second, more complex dataset, the optimized eigenvalues referred to the first layer are a good proxy of the importance of the associated components.
@@ 3.3 The MNIST dataset
As an additional test, we applied the spectral features detection method to a reduced version of the well-known MNIST dataset [29]. This latter consists of $28\times 28$ images depicting handwritten digits, ranging from 00 to $9$, to be eventually classified. In our experiments, and to deliver a clear message, we first focused on the subset of images “zeroes" and “ones". Despite the simplicity of the adopted framework, it is interesting to test the method against a real dataset, where the measure of the relevance of the components is not a priori given. The input vectors of the MNIST dataset has a dimension of $784$, and each component of the vectors represents the intensity of a specific pixel in the corresponding image.
@@
Once again, we trained a spectral feedforward neural network by adjusting the trainable spectral parameters. The hidden structure of the network is identical to that employed for the above applications. Clearly, the input layer is now made of 784 elements, the size of the supplied items. The exit layer is still composed of just two nodes, one associated to the images that display a zero and the other pointing to the ones. As expected, the trained model performs very well in terms of estimated classification accuracy (larger than $99$).
@@
Finally, we also report in Figure 9 the recorded accuracy for the complete collection of digits (ranging from 00 to $9$) which compose the entire MNIST dataset. Also in this case, and in analogy with the results obtained for the examined settings in which pair of digits are directly compared, the accuracy raises as a function of the number of input pixels included in the analysis, and ranked according to their spectral importance.
@@ 3.4 Stellar spectra and metallicity prediction
In this Section we report on the performance of the algorithm against a dataset of interest for astrophysics. More specifically,we consider the MILES stellar library [30], composed of 985 stars. The dataset is composed of stellar spectra, namely flux as a function of wavelength, in the range 3525-7500 AA. For each spectrum we consider the associated stellar parameters (effective temperature, gravity, and the abundance of heavy elements known as metallicity) from [31]. A few representative spectra are shown in Figure 10. Because of its spectral type coverage and flux calibration accuracy this stellar library has been extensively used as a basis for constructing stellar population models to analyze galaxy spectra [32, 33] and therefore deriving metallicity and star-formation histories of galaxies.
Figure 10: Plot of some of the stellar spectra composing the MILES dataset [30].
@@
Figure 11: In the upper-left panel the normalized distribution of the eigenvalues of the input layer for the metallicity prediction task is reported in vertical log scale. Upper-right panel shows the mean absolute error of the model versus the fraction of spectral lines provided as an input. The order in which the lines are inserted follows the absolute value (magnitude) of their corresponding post-training eigenvalues. The bottom panel shows the wavelength ranges in the stellar spectra which are deemed relevant for the metallicity prediction task (grey shadows) along with the first layer eigenvalues (see also Method section)
We chose metallicity as the target because it can be inferred from well-defined absorption lines of metal species. Such lines are, however, generally blended at the resolution of our dataset and in galaxy spectra. A set of “indices”, or wavelength regions, have therefore been defined [34, 35], which are sensitive to various combinations of metallicity, temperature, and gravity. We will compare the feature importances obtained from the machine learning analysis with these classical indices.
@@
The upper-left panel in Figure (11) shows the normalized eigenvalues distribution of the input layer eigenvalues. As for the other examples discussed above, a clear peak is displayed close to zero (note the logarithmic scale on the $y$-axis) thus implying that several input features play no role for the handled task. In particular, if we compute the prediction error on the test set as a function of the number of spectral lines used as input — ordered according to their corresponding eigenvalues — we observe that the minimum error is reached using only around $30$ lines (see the upper-right panel in Figure 11). In other words, the network is able to make accurate predictions by relying on just $0.7\%$ of the input components. The eigenvalues allow hence to identify a subset of relevant input elements that drive decision making across the network. In this respect, the algorithm is thus successfully tested against a real dataset with non trivial physical implications.

### v3

@@ 2.2 Correlated Gaussian distributions dataset
$\begin{split}&x_{2n}\sim\mathcal{N}(0,1),\\ &x_{2n+1}=x_{2n}\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\text{if}\;\;c(\vec{x})=0,\\ &x_{2n+1}=\begin{cases}x_{2n}\;\text{with prob.}\;1-p_{n}\\ -x_{2n}\;\text{with prob.}\;p_{n}\end{cases}\;\;\text{if}\;\;c(\vec{x})=1,\end{split}$ (2)
where $c(\vec{x})$ is the target class of the input and the index $n\in\{0,1,2,3\}$ denotes the four pairs. The last two components are independently drawn from two identical normal distributions $\mathcal{N}(0,1)$ regardless of the class of belonging:
@@
To better understand the relationship between the different components, let us focus on the first two entries of the input vectors, $x_{0}$ and $x_{1}$. For data belonging to class 0 (as we label the first of the two) $x_{0}=x_{1}$, that is, the first two components are perfectly correlated. Conversely, the two components of data belonging to the other class (denoted by $1$) are equal with probability $1-p_{0}$ and have opposite sign with probability $p_{0}$. If $p_{0}=0$, the first two components are equal for both classes. No information about the class to which the supplied data belong can be thus extracted by comparing the two above components. The second component is indeed a mere copy of the first and therefore irrelevant for classification purposes. On the other hand, if $p_{0}=1$, the first and second components from class $1$ data are in perfect anti-correlation, while data from class 0 are still matched and thus fully correlated. Hence, in the latter case, it is possible to completely separate the two classes by solely comparing $x_{0}$ and $x_{1}$. The Parameter $p_{n}$ quantifies therefore the degree of sensible information which can be pulled out from a direct comparison of the two components in the $n$-th pair. We want to stress that, regardless of the value of $p_{n}$, neither component individually contains information on the specific class to which the vector belongs. Therefore, useful information can be obtained to solve the classification task only by analyzing the two components simultaneously. Following this observations, we generated the dataset by using four different values of $p_{n}\in[0,1]$ for the four pairs. Finally, the last two components are not correlated with the others and are identically distributed for both classes, hence void of information. We therefore assign a parameter $p_{4}=0$ to these two components to explicitly state their manifest irrelevance. Note that the two classes are not linearly separable, even in the case one couple of components is associated to a $p_{n}=1$. A total of $10^{5}$ examples were generated to perform the test.
@@
The architecture of the employed neural network, parameterized via its spectral modality, is identical to that put forward for the analysis of the uncorrelated dataset (except for the first layer that in this case receives just 10 entries). In upper right panel of Fig. 4, the post-training eigenvalues referred to the input layer are shown. Firstly, we notice that pairs are correctly spotted out, as they are assigned the very same eigenvalue entry. Moreover, eigenvalues associated with irrelevant components are tuned zero for this dataset as well, an observation that testifies again on the ability of the spectral network to discriminate between irrelevant and relevant features. Lastly, the magnitude of the recorded eigenvalues scales proportionally to the control parameter $p_{n}$, which quantifies the relevance of the selected pair for classification purposes (bottom right panel, Fig. 4). The Pearson correlation ($\rho$) between post-training eigenvalues and the relative $p_{n}$ values was measured over ten independent training runs, yielding $\rho=0.988\pm 0.004$. In conclusion, and also for this second, more complex dataset, the optimized eigenvalues referred to the first layer are a good proxy of the importance of the associated components.
@@ 2.3 The MNIST dataset
As an additional test, we applied the spectral features detection method to a reduced version of the well-known MNIST dataset [deng2012mnist]. This latter consists of $28\times 28$ images depicting handwritten digits, ranging from 0 to $9$, to be eventually classified. In our experiments, and to deliver a clear message, we first focused on the subset of images “zeroes" and “ones". Despite the simplicity of the adopted framework, it is interesting to test the method against a real dataset, where the measure of the relevance of the components is not a priori given. The input vectors of the MNIST dataset has a dimension of $784$, and each component of the vectors represents the intensity of a specific pixel in the corresponding image.
@@
Once again, we trained a spectral feedforward neural network by adjusting the trainable spectral parameters. The hidden structure of the network is identical to that employed for the above applications. Clearly, the input layer is now made of 784 elements, the size of the supplied items. The exit layer is still composed of just two nodes, one associated to the images that display a zero and the other pointing to the ones. As expected, the trained model performs very well in terms of estimated classification accuracy (larger than $99$ %).
@@
Finally, we also report in Figure 9 the recorded accuracy for the complete collection of digits (ranging from 0 to $9$) which compose the entire MNIST dataset. Also in this case, and in analogy with the results obtained for the examined settings in which pair of digits are directly compared, the accuracy raises as a function of the number of input pixels included in the analysis, and ranked according to their spectral importance.
@@ 2.4 Stellar spectra and metallicity prediction
In this Section we report on the performance of the algorithm against a dataset of interest for astrophysics. More specifically, we consider the MILES stellar library [sanchez2006medium], composed of 985 stars. The dataset is composed of stellar spectra, namely flux as a function of wavelength, in the range 3525-7500 Å. For each spectrum we consider the associated stellar parameters (effective temperature, gravity, and the abundance of heavy elements known as metallicity) from [cenarro2007medium]. A few representative spectra are shown in Figure 10. Because of its spectral type coverage and flux calibration accuracy this stellar library has been extensively used as a basis for constructing stellar population models to analyze galaxy spectra [vazdekis2010evolutionary, vazdekis2016uv] and therefore deriving metallicity and star-formation histories of galaxies.
Figure 10: Plot of some of the stellar spectra composing the MILES dataset [sanchez2006medium].
@@
Figure 11: In the upper-left panel the normalized distribution of the eigenvalues of the input layer for the metallicity prediction task is reported in vertical log scale. Upper-right panel shows the mean absolute error of the model versus the fraction of spectral lines provided as an input. The order in which the lines are inserted follows the absolute value (magnitude) of their corresponding post-training eigenvalues. The dashed line refers to the absolute error made by a standard MLP with the same architecture. The bottom panel shows the wavelength ranges in the stellar spectra which are deemed relevant for the metallicity prediction task (grey shadows) along with the first layer eigenvalues (see also Method section).
We chose metallicity as the target because it can be inferred from well-defined absorption lines of metal species. Such lines are, however, generally blended at the resolution of our dataset and in galaxy spectra. A set of “indices”, or wavelength regions, have therefore been defined [faber1985old, worthey1994old], which are sensitive to various combinations of metallicity, temperature, and gravity. We will compare the feature importances obtained from the machine learning analysis with these classical indices.
@@
The upper-left panel in Figure (11) shows the normalized eigenvalues distribution of the input layer eigenvalues. As for the other examples discussed above, a clear peak is displayed close to zero (note the logarithmic scale on the $y$-axis) thus implying that several input features play no role for the handled task. In particular, if we compute the prediction error on the test set as a function of the number of spectral lines used as input — ordered according to their corresponding eigenvalues — we observe that the minimum error is reached using only around $30$ lines (see the upper-right panel in Figure 11). In other words, the network is able to make accurate predictions by relying on just $0.7\%$ of the input components. Notably, the measured error is comparable to that of a standard MLP (see top right panel in figure 11). The eigenvalues allow hence to identify a subset of relevant input elements that drive decision making across the network. In this respect, the algorithm is thus successfully tested against a real dataset with non trivial physical implications.
3 Comparison with other methods
In this section, we briefly compare the results obtained using our proposed method with those obtained by employing other approaches, as available in the literature, to carry out the same task. This comparison is not intended to be exhaustive, and a more in-depth analysis between the spectral method in relation to other techniques is left for future work.
SHapley Additive exPlanations (SHAP). A first comparison was made with the SHAP method [lundberg2017unified], a well-known framework based on game theory which is able to explain how input features contribute to the model’s prediction. Since our method provides a measure of the global relevance of input components, independent of any specific item being inspected, we derived comparable information by averaging the importance values obtained by using SHAP over multiple data points. First of all, we trained a model using the spectral procedure on the simplified MNIST dataset, as described in section 2.3. We then compared the normalized eigenvalues (i.e., the eigenvalues multiplied by the norm of the corresponding eigenvectors) as obtained at the end of the training with the average pixel importance values provided by SHAP on the same model. In the upper panels of Figure (12), the relevance values estimated by using the two methods are shown in the input space. The results are very similar: the pixels identified as most relevant are distributed in the same way within the image. In the bottom-left panel, the normalized eigenvalues are plotted against the SHAP-based relevance scores, showing good agreement between the two estimators. This confirms that post-training eigenvalues provide reliable estimates of the input component importance, comparable to the relevance scores produced by the SHAP approach. However, since eigenvalues are parameters tuned during the training phase and are inherently input-independent, the computational cost of the spectral method is significantly lower. To prove this, in the bottom-right panel of Figure (12), the computation time required to obtain SHAP relevance scores is reported for varying sizes of the portion of the dataset used for averaging. As the subset size increases, the computation time scales linearly. Nevertheless, even for very small subsets, the SHAP computation time is orders of magnitude higher than the time required to normalize the eigenvalues of the first layer, which is the only post training operation required when dealing with the spectral method.
It is important to note that methods such as SHAP are limited to providing an estimate of the importance of input features, but do not include strategies for guiding models toward solutions that rely on few input components. Such strategies must be formulated separately. In fact, when examining the scores obtained using the SHAP approach on a standard (non-spectral) model with the same architecture and hyperparameters, it can be observed that the scores are nonzero across most of the input domain, even when a regularization term is applied to the weights of the first layer (Figure 13, left). In contrast, using the spectral formulation, it is possible to guide the model during training toward solutions that are as sparse as possible in terms of the input components used. This is achieved by introducing a regularization term on the eigenvalues of the input layer. To emphasize this, we apply an $L_{1}$ regularization on the eigenvalues, enforcing solutions that are even sparser than those discussed earlier in this section and in Section (2.3) (Figure 13, right).
Figure 12: The two upper panels show the importance values obtained by using (i) the SHAP method normalized by their maximum (left) and (ii) the normalized eigenvalues (right). The results refer to a model trained using the spectral procedure on the simplified MNIST dataset (see Section 2.3), with values represented in the input space. The bottom-left panel plots the normalized eigenvalues against the importance values obtained using the SHAP method. In the bottom-right panel, the computation times required to calculate the normalized eigenvalues (green dashed line) and SHAP values (orange diamonds) are shown for different sizes of the dataset used to estimate the average SHAP scores.
Figure 13: The left panel shows the importance values obtained using the SHAP method normalized by their maximum (left) and for a standard MLP with a $L_{1}$ regularization factor on the weights of the first layer. The right panel shows the normalized eigenvalues of the first layer of a spectral neural network where a $L_{1}$ regularization factor has been used on the eigenvalues. Both the results refer to the simplified MNIST and the values are plotted in the input domain.
Layer-Wise relevance propagation. The second technique against which we compare our proposed method is Layer-wise Relevance Propagation (LRP) [montavon2019layer]. More specifically, we do not present here a comparison of the respective experimental results. Instead, we elaborate on the theoretical framework underlying the LRP method and explain how it can be naturally reconciled with the spectral parameterization. Let us begin by recalling the basic aspects that underlie the LRP method. The LRP technique aims to estimate the relevance $R^{l}_{i}$ of nodes by propagating relevance scores backward from the output to the input. The relevance $R^{l}_{i}$ of a node in layer $l$ is related to the relevance $R^{l+1}_{j}$ of the nodes in layer $l+1$ through the linear relationship:
$R^{l}_{i}=\sum_{j}\frac{z_{ij}}{\sum_{k}z_{kj}}R^{l+1}_{j}.$ (4)
The elements $z_{ij}$ define how the relevance scores are distributed when they are back-propagated from a layer to the previous one. Different choices for such elements and for the employed normalization term appearing in the denominator yield different rules for LRP. Some common rules are:
$\text{LRP-0 rule:}\hskip 56.9055ptR^{l}_{i}=\sum_{j}\frac{a_{i}w_{ij}}{\sum_{k}a_{k}w_{kj}}R^{l+1}_{j},$
$\text{LRP-$\epsilon$ rule:}\hskip 56.9055ptR^{l}_{i}=\sum_{j}\frac{a_{i}w_{ij}}{\epsilon+\sum_{k}a_{k}w_{kj}}R^{l+1}_{j},$
 $\text{ LRP-$\gamma$ rule:}\hskip 56.9055ptR^{l}_{i}=\sum_{j}\frac{a_{i}(w_{ij}+\gamma w_{ij}^{+})}{\sum_{k}a_{k}(w_{kj}+\gamma w_{kj}^{+})}R^{l+1}_{j},$
where $w_{ij}$ is the weight of the connection between neurons $i$ and $j$, and $a_{i}$ is the activation observed at neuron $i$ when the input data is provided to the model. This clearly implies that the method is input-dependent. Furthermore, different LRP rules can be derived assuming a first order Taylor expansion of the Deep Neural Network around an Input Dependent point, therefore specifically intertwining the value of the back-propagated relevance and the input.
These rules can be rewritten under the assumption of a spectral parameterization by substituting $w_{ij}\rightarrow\lambda_{i}\psi_{ij}$. Focusing on the LRP-0 rule, we have:
$R^{l}_{i}=\sum_{j}\frac{a_{i}\lambda_{i}\psi_{ij}}{\sum_{k}a_{i}\lambda_{k}\psi_{kj}}R^{l+1}_{j}=\lambda_{i}\sum_{j}\frac{a_{i}\psi_{ij}}{\sum_{k}a_{i}\lambda_{k}\psi_{kj}}R^{l+1}_{j},$ (5)
from which it is clear that the relevance $R^{l}_{i}$ of node $i$ is modulated by the corresponding eigenvalue $\lambda_{i}$. This result can be easily extended to $\epsilon$-rule. Moreover, by assuming $\lambda_{i}>0$ $\forall i$, the result holds also for the $\gamma$-rule, in fact:
$a_{i}(w_{ij}+\gamma w_{ij}^{+})=a_{i}(\lambda_{i}\psi_{ij}+\gamma\lambda_{i}\psi_{ij}^{+})\propto\lambda_{i}.$ (6)
The condition $\lambda_{i}>0$ is typically satisfied since the eigenvalues are initialized to one (see Methods), but it can also be enforced as a constraint during training. Finally, the global relevance of a node (i.e., input-independent) provided by the LRP method can be computed by averaging over multiple data points. However, the eigenvalues are not affected by this averaging process, and the global relevance of a node remains modulated by its corresponding eigenvalue, namely:
$\langle R^{l}_{i}\rangle\propto\lambda_{i},$ (7)
where $\langle R^{l}_{i}\rangle$ denotes the average relevance of node $i$ across the dataset. A complete analysis of the consequences of this relation is left for future work. Once again, as for the SHAP method, it is important to emphasize that LRP is an input-dependent method and requires extensive post-training computations to estimate relevance. Moreover, it does not provide any mechanism for guiding the model toward solutions that rely on a minimal set of input components, as it is instead possible via the spectral paradigm complemented with suitable regularization terms.
Gradient-based methods. Finally, we compare our method with some gradient-based techniques. In particular, we considered two distinct approaches: the first is essentially a vanilla saliency map, where the importance of the input features is directly associated with the gradient of the output with respect to the components of the input vector. The second is a more advanced technique known as Integrated Gradients [sundararajan2017axiomatic], which we here choose to apply by setting a zero vector as the baseline. Since both methods provide input-dependent relevance estimates, we averaged the results over multiple elements to obtain a global relevance estimate.
Figure 14: Wavelength ranges in the stellar spectra which are deemed relevant for the metallicity prediction task (grey shadows) along with the first layer eigenvalues (orange diamonds), saliency values (green stars) and importance values given by integrate gradients (blue triangles). Only the values larger than $10^{-2}$ are reported.
In Figure (14), we report the results obtained for the stellar spectra dataset, as introduced above. The model was trained using the spectral formulation with $L_{1}$ regularization applied to the eigenvalues and eigenvectors of the input layer. The three methods (spectral, saliency map, and integrated gradients) identify the same wavelengths as essential for solving the regression task, assigning comparable importance scores to those wavelengths. Indeed, the correlation between the post-training eigenvalues and the saliency map scores for this task (computed over ten training runs of the model) is $0.98\pm 0.01$, while the correlation with the integrated gradients scores is $0.95\pm 0.02$. We have also measured the computation time required to obtain the global score using the three different methods during the same training runs. The spectral method proved to be the fastest, with a computation time of $t=0.003\text{s}\pm 0.0005\text{s}$. Since an average over multiple samples is required to compute the other two global (input-independent) scores, their computation times depend on the number of samples over which the average is performed. Even when averaging over only $20$ samples (the size of the test set for this task), the computation time for the global saliency map scores is $t=0.34\text{s}\pm 0.03\text{s}$, while for the integrated gradients it is $t=0.58\text{s}\pm 0.02\text{s}$. A more exhaustive analysis of the computational time and algorithmic complexity is reported in Appendices (C) and (D).). In any cases, the observed reduction in the number of relevant components is a consequence of applying regularization during training to the new parameters introduced by the spectral formulation, namely the eigenvalues, a possibility that is not available when working with a standard MLP.
From all these comparisons, we can conclude that the post-training eigenvalues of the first layer provide a sensible ranking of the input components, in line with that recovered with other available techniques. This demonstrates that the eigenvalues serve as a reliable proxy for the global (i.e., input-independent) relevance of the input components. Moreover, we showed how the use of spectral formulation is an intrinsically different approach compared to the methods described above. In fact, it not only provides a means to estimate relevance, but also offers a strategy to easily enforce solutions that rely on sparse representations of the data. Let us sum up some pros and cons of the proposed method:
Pros:
  • Input-independent method. The eigenvalues serve as an effective proxy for the relevance of the input components across the entire dataset. Furthermore, no additional computation is required to aggregate these relevance measures over multiple data instances.
  • Very fast algorithm. The only computation required to obtain the relevance on the input components is to normalize the first-layer eigenvalues (i. e., to multiply them by the norm of the associated eigenvectors).
  • Enforce sparse solutions. As the parameters approximating feature relevance are trainable components of the model, they can be regularized to favour solutions that rely on a minimal set of input components.
Cons:
  • Input-independent method. Having an input-independent method can be disadvantageous when one wants to understand the decision-making process that led to a prediction for a specific input. A possible solution is to overlay the global relevance map onto the specific input to extract data-dependent importance values, as reported in Figure 7.
  • Model-dependent. The method applies solely to neural network models formulated with spectral parameterization.